can y'all help with with these 3 please!! Thank you!!

Question 1

Find the volume beneath z = f(x,y) and above the region described by the triangle with vertices (0,0), (4,0), and (0,4).

f(x,y)= -x-y+c ; use c = 7.

Hint: compute the double integral required to find the volume under f(x,y) using the limits of integration given by the region on the x-y plane.

Question 2

Prove that F is a gradient field and determine the work of F to go from point A: (1 , 3, 2) to point B: (4 , 2 , 4) along any trajectory.

F=[2xyz x2z+2e(-2y) x2y]

Question 5

Determine the line integral along the curve C from A to B. Find the parametric form of the curve C. Use the vector field:



Use the following values: a=5; b=2; and c=6.
Answer #1

4 Som & 2 - flasy) = x-y+7 v-ff e dady- y) andy hty+z=7 x+y+z=7 12 2 SE*-**-yagedy 1128-8-49 Joly 1120-]dx - (204-29 27-80-20